# SDEs with supercritical distributional drifts

### v1

@@ 1. $\bullet$
Let
$u_{n}(t,x):=u(t,\cdot)*\phi_{n}(x),$
where $\phi_{n}(x)=n^{d}\phi(nx)$ with $\phi\in C^{\infty}_{c}({\mathbb{R}}^{d})$ is a smooth probability density function. Then by Remark 3.4 with $\varphi=\phi_{n}(x-\cdot)$, we get
$\partial_{s}u_{n}+\Delta u_{n}-\lambda u_{n}=-\phi_{n}*(b\odot\nabla u-\mathord{{\rm div}}b\preceq u+b)=:g_{n},$ (4.20)
@@
Observe that by definition,
$X_{t}=X_{0}+A^{b}_{t}+\sqrt{2}W_{t},$
where $A^{b}_{t}:=\lim_{n\to\infty}\int^{t}_{0}b_{n}(s,X_{s}){\mathord{{\rm d}}}s$ by Proposition 4.3, satisfies
$\sup_{t\in[0,T]}\|A^{b}_{t}\|_{L^{m}(\Omega)}+\sup_{s\neq t\in[0,T]}\frac{\|A^{b}_{t}-A^{b}_{s}\|_{L^{m}(\Omega)}}{|t-s|^{(1+\theta_{0})/2}}\lesssim_{C}\|b\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}.$
@@
$Y^{n}_{t}$ $:=\Phi_{n}(t,X_{t})=\Phi_{n}(0,x)+\int^{t}_{0}(\partial_{s}u_{n}+\Delta u_{n})(s,X_{s}){\mathord{{\rm d}}}s$
$\qquad+\int^{t}_{0}{\mathord{{\rm d}}}A^{b}_{s}\cdot\nabla\Phi_{n}(s,X_{s})+\sqrt{2}\int^{t}_{0}\nabla\Phi_{n}(s,X_{s}){\mathord{{\rm d}}}W_{s}$
$=\Phi_{n}(0,x)+\lambda\int^{t}_{0}u_{n}(s,X_{s}){\mathord{{\rm d}}}s+A^{b\cdot\nabla\Phi_{n}+g_{n}}_{t}+\sqrt{2}\int^{t}_{0}\nabla\Phi_{n}(s,X_{s}){\mathord{{\rm d}}}W_{s}.$ (4.21)
@@
$b\cdot\nabla\Phi_{n}+g_{n}$ $=b+b\cdot\nabla u_{n}-\phi_{n}*(b\odot\nabla u-\mathord{{\rm div}}b\preceq u+b)=h^{n}_{1}+h^{n}_{2},$
@@
$h^{n}_{1}:=b-b*\phi_{n}+b\odot\nabla u_{n}-\phi_{n}*(b\odot\nabla u),\ \ h^{n}_{2}:=\mathord{{\rm div}}b\preceq u_{n}-\phi_{n}*(\mathord{{\rm div}}b\preceq u).$
@@
${\mathbf{E}}|A^{h^{n}_{1}}_{t}|\lesssim\|h^{n}_{1}\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}$ $\lesssim\|b-b*\phi_{n}\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}+\|b\odot\nabla(u_{n}-u)\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}$
$\quad+\|\phi_{n}*(b\odot\nabla u)-b\odot\nabla u\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}.$
@@
$\|b\odot\nabla(u_{n}-u)\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}$ $\lesssim\kappa^{b}_{1}\|u_{n}-u\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{1+d/p_{b}}_{p_{b},1}}\lesssim n^{-\varepsilon}.$
@@
so that $(-\varepsilon,p_{b}/2,q_{b}/2)\in{\mathscr{I}}^{\alpha_{b}}_{p_{b},q_{b}}$. By Krylov’s estimate again, we have
${\mathbf{E}}|A^{h^{n}_{2}}_{t}|\lesssim\|h^{n}_{2}\|_{{\mathbb{L}}^{q_{b}/2}_{T}{\mathbf{B}}^{-\varepsilon}_{p_{b},q_{b}}}$ $\lesssim\|\phi_{n}*(\mathord{{\rm div}}b\preceq u)-\mathord{{\rm div}}b\preceq u\|_{{\mathbb{L}}^{q_{b}/2}_{T}{\mathbf{B}}^{-\varepsilon}_{p_{b}/2,q_{b}}}$
$+\|\mathord{{\rm div}}b\preceq(u_{n}-u)\|_{{\mathbb{L}}^{q_{b}/2}_{T}{\mathbf{B}}^{-\varepsilon}_{p_{b},q_{b}}}=:I_{n}+J_{n}.$
Note that by (2.17),
$\lim_{n\to\infty}J_{n}$ $\lesssim\lim_{n\to\infty}\|\mathord{{\rm div}}b\preceq(u_{n}-u)\|_{{\mathbb{L}}^{q_{b}/2}_{T}{\mathbf{B}}^{0}_{p_{b},\infty}}$
$\lesssim\kappa^{b}_{2}\lim_{n\to\infty}\|u_{n}-u\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{2+\alpha_{b}}_{p_{b},q_{b}}}=0.$
Since $\|\mathord{{\rm div}}b\preceq u\|_{{\mathbb{L}}^{q_{b}/2}_{T}{\mathbf{B}}^{-\varepsilon}_{p_{b}/2,q_{b}}}\lesssim\|\mathord{{\rm div}}b\preceq u\|_{{\mathbb{L}}^{q_{b}/2}_{T}{\mathbf{B}}^{0}_{p_{b}/2,\infty}}<\infty$, by the property of convolution, we also have
@@
$\lim_{n\to\infty}{\mathbf{E}}|A^{b\cdot\nabla\Phi_{n}+g_{n}}_{t}|=0.$
@@
[…]
$C_{0}t^{-d/2}\mathrm{e}^{-c_{0}|x-x^{\prime}|^{2}/t}\leqslant\rho_{t}(x,x^{\prime})\leqslant C_{1}t^{-d/2}\mathrm{e}^{-c_{1}|x-x^{\prime}|^{2}/t},$ (4.22)
where $C_{0},C_{1},c_{0},c_{1}>0$ only depend on the parameters $T,d,\alpha_{b},p_{b},q_{b}$ and $\|b\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}$.
@@
(Existence). Consider the approximation SDE (4.11). By Lemma 4.5, the law of $X^{n}$ in ${\mathbb{C}}_{T}$ is tight, and therefore, by Prohorov’s theorem, it is relatively compact in ${\mathcal{P}}({\mathbb{C}}_{T})$. Then, by the weak convergence method, it is standard to derive the existence of a weak solution (see [47] and also the proof of Theorem 5.7 below).
(Uniqueness and heat kernel estimate). By Lemma 4.7, since the transformed SDE (4.17) has a bounded Lipschitz drift and bounded uniformly non-degenerate Hölder diffusion coefficients, it is well known that SDE (4.17) has a unique weak solution $Y_{t}$ (see [38]), and for each $t\in(0,T]$ and starting point $Y_{0}=y$, $Y_{t}$ has a density $\tilde{\rho}_{t}(y,y^{\prime})$, which enjoys two sides Gaussian estimate (for example, see [10]):
@@
The desired estimate (4.22) follows by the change of variables (see [42]). ∎
@@
Since $b$ is divergence free and by Lemma 2.4,
@@
[…]

### v3

@@ 3.2. Supercritical case: $\frac{d}{p}+\frac{2}{q}<2+\alpha$
Let $\phi_{n}(x)$ be the mollifier as in (1.2). Define
$u_{n}(t,x):=u(t,\cdot)*\phi_{n}(x).$
Then by Remark 3.4 with $\varphi=\phi_{n}(x-\cdot)$, we get
$\partial_{s}u_{n}+\Delta u_{n}-\lambda u_{n}=-\phi_{n}*({b_{1}}\odot\nabla u-\mathord{{\rm div}}{b_{1}}\preceq u+b)=:g_{n},$ (4.19)
@@
By Krylov’s estimate and definition, it is easy to see that
$X_{t}=X_{0}+A^{b}_{t}+\sqrt{2}W_{t}=X_{0}+A^{b_{1}}_{t}+\int^{t}_{0}b_{2}(s,X_{s}){\mathord{{\rm d}}}s+\sqrt{2}W_{t},$
where $A^{b_{1}}_{t}:=\lim_{n\to\infty}\int^{t}_{0}b_{1,n}(s,X_{s}){\mathord{{\rm d}}}s$ by Proposition 4.3, satisfies
$\sup_{t\in[0,T]}\|A^{b_{1}}_{t}\|_{L^{m}(\Omega)}+\sup_{s\neq t\in[0,T]}\frac{\|A^{b_{1}}_{t}-A^{b_{1}}_{s}\|_{L^{m}(\Omega)}}{|t-s|^{(1+\theta_{0})/2}}\leqslant C\|{b_{1}}\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}.$
@@
$Y^{n}_{t}$ $:=\Phi_{n}(t,X_{t})=\Phi_{n}(0,x)+\int^{t}_{0}(\partial_{s}u_{n}+\Delta u_{n})(s,X_{s}){\mathord{{\rm d}}}s+\int_{0}^{t}(b_{2}\cdot\nabla u_{n})(s,X_{s}){\mathord{{\rm d}}}s$
$\qquad+\int^{t}_{0}{\mathord{{\rm d}}}A^{b_{1}}_{s}\cdot\nabla\Phi_{n}(s,X_{s})+\sqrt{2}\int^{t}_{0}\nabla\Phi_{n}(s,X_{s}){\mathord{{\rm d}}}W_{s}$
$=\Phi_{n}(0,x)+\lambda\int^{t}_{0}u_{n}(s,X_{s}){\mathord{{\rm d}}}s+A^{b_{1}\cdot\nabla\Phi_{n}+g_{n}}_{t}+\int_{0}^{t}(b_{2}\cdot\nabla u_{n})(s,X_{s}){\mathord{{\rm d}}}s$
$\qquad+\sqrt{2}\int^{t}_{0}\nabla\Phi_{n}(s,X_{s}){\mathord{{\rm d}}}W_{s}.$ (4.20)
@@
${b_{1}}\cdot\nabla\Phi_{n}+g_{n}$ $={b_{1}}+{b_{1}}\cdot\nabla u_{n}-\phi_{n}*({b_{1}}\odot\nabla u-\mathord{{\rm div}}{b_{1}}\preceq u+{b_{1}})=h^{n}_{1}+h^{n}_{2},$
@@
$h^{n}_{1}:={b_{1}}-{b_{1}}*\phi_{n}+{b_{1}}\odot\nabla u_{n}-\phi_{n}*({b_{1}}\odot\nabla u),\ \ h^{n}_{2}:=\mathord{{\rm div}}{b_{1}}\preceq u_{n}-\phi_{n}*(\mathord{{\rm div}}{b_{1}}b\preceq u).$
@@
${\mathbf{E}}|A^{h^{n}_{1}}_{t}|\lesssim\|h^{n}_{1}\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}$ $\lesssim\|{b_{1}}-{b_{1}}*\phi_{n}\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}+\|{b_{1}}\odot\nabla(u_{n}-u)\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}$
$\quad+\|\phi_{n}*({b_{1}}\odot\nabla u)-{b_{1}}\odot\nabla u\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}.$
@@
$\|{b_{1}}\odot\nabla(u_{n}-u)\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}$ $\lesssim\kappa_{b}\|u_{n}-u\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{1+d/p_{b}}_{p_{b},1}}\lesssim n^{-\varepsilon}.$
@@
so that $(-\varepsilon,p_{b}/2,q_{b}/2)\in{\mathbf{I}}^{\alpha_{b}}_{p_{b},q_{b}}$. By Krylov’s estimate again, we have
${\mathbf{E}}|A^{h^{n}_{2}}_{t}|\lesssim\|h^{n}_{2}\|_{{\mathbb{L}}^{q_{b}/2}_{T}{\mathbf{B}}^{-\varepsilon}_{p_{b},q_{b}}}$ $\lesssim\|\phi_{n}*(\mathord{{\rm div}}{b_{1}}\preceq u)-\mathord{{\rm div}}{b_{1}}\preceq u\|_{{\mathbb{L}}^{q_{b}/2}_{T}{\mathbf{B}}^{-\varepsilon}_{p_{b}/2,q_{b}}}$
$+\|\mathord{{\rm div}}{b_{1}}\preceq(u_{n}-u)\|_{{\mathbb{L}}^{q_{b}/2}_{T}{\mathbf{B}}^{-\varepsilon}_{p_{b},q_{b}}}=:I_{n}+J_{n}.$
For $J_{n}$, by (2.17) we have
$\lim_{n\to\infty}J_{n}\lesssim\lim_{n\to\infty}\|\mathord{{\rm div}}{b_{1}}\preceq(u_{n}-u)\|_{{\mathbb{L}}^{q_{b}/2}_{T}{\mathbf{B}}^{0}_{p_{b},\infty}}\lesssim\kappa^{b_{1}}_{2}\lim_{n\to\infty}\|u_{n}-u\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{2+\alpha_{b}}_{p_{b},q_{b}}}=0.$
For $I_{n}$, since $\|\mathord{{\rm div}}{b_{1}}\preceq u\|_{{\mathbb{L}}^{q_{b}/2}_{T}{\mathbf{B}}^{-\varepsilon}_{p_{b}/2,q_{b}}}\lesssim\|\mathord{{\rm div}}{b_{1}}\preceq u\|_{{\mathbb{L}}^{q_{b}/2}_{T}{\mathbf{B}}^{0}_{p_{b}/2,\infty}}<\infty$, by the property of convolution, we also have
@@
$\lim_{n\to\infty}{\mathbf{E}}|A^{{b_{1}}\cdot\nabla\Phi_{n}+g_{n}}_{t}|=0.$
@@
To show the existence of a weak, we consider the following approximation SDE:
${\mathord{{\rm d}}}X^{n}_{t}=b_{n}(t,X^{n}_{t}){\mathord{{\rm d}}}t+\sqrt{2}{\mathord{{\rm d}}}W_{t},\ \ X^{n}_{0}\sim\mu\in{\mathcal{P}}({\mathbb{R}}^{d}),$ (4.21)
where $b_{n}:=b_{1,n}+b_{2}$ and $b_{1,n}(t,x):=b_{1}(t,\cdot)*\phi_{n}(x)$. Since $\int^{T}_{0}\|b_{1,n}(t,\cdot)\|^{q_{b}}_{\infty}{\mathord{{\rm d}}}t<\infty$ and $|b_{2}(t,x)|\leqslant c_{0}+c_{1}|x|$, it is well known that the weak well-posedness holds for SDE (4.21) (see [50]). Note that by the property of convolution,
$\begin{split}\kappa_{b_{1,n}}:=\|b_{1,n}\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}&\leqslant\|b_{1}\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}=\kappa_{b_{1}},\\ \tilde{\kappa}_{b_{1,n}}:=\|\mathord{{\rm div}}b_{1,n}\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{-2-\alpha_{b}}_{p_{b},q_{b}/(q_{b}-1)}}&\leqslant\|\mathord{{\rm div}}b_{1}\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{-2-\alpha_{b}}_{p_{b},q_{b}/(q_{b}-1)}}=\tilde{\kappa}_{b_{1}}.\end{split}$ (4.22)
Now we can show the following crucial Krylov estimate.
Lemma 4.6.
Let $\mu\in{\mathcal{P}}({\mathbb{R}}^{d})$ have finite $m$-order moment, where $m$ satisfies (1.5). For any $(\alpha,p,q)\in{\mathbf{I}}^{\alpha_{b}}_{p_{b},q_{b}}$, $(X^{n},{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q})$ satisfies Krylov’s estimate (4.3) with parameters $(m,\theta_{0})$ and constant $C=C(\Theta,\alpha,p,q,m,\mu,c_{0},c_{1})$, where $c_{0},c_{1}$ are from ($\widetilde{\bf H}^{\rm sub}_{b}$), and $C=C(\Theta,\alpha,p,q,c_{0})$ when $c_{1}=0$.
Proof.
We divide the proof into two steps.
(Step 1). We establish the following uniform moment estimate: for some $C_{m}=C_{m}(\Theta,c_{0},c_{1})>0$,
$\sup_{n}\left\|\sup_{t\in[0,T]}|X^{n}_{t}|\right\|_{L^{m}(\Omega)}\leqslant C_{m}\big{(}1+(\mu(|\cdot|^{m}))^{1/m}\big{)}=:\ell_{m}.$ (4.23)
Let $\Phi_{n}$ be defined as in (4.14), where $b_{1,n}$ is used in place of $b_{1}$. Define $Y^{n}_{t}:=\Phi_{n}(t,X^{n}_{t})$. Using SDE (4.16), BDG’s inequality, (4.17), and (4.18), we obtain:
$\left\|\sup_{t\in[0,T]}|Y^{n}_{t}|\right\|_{L^{m}(\Omega)}\lesssim\|\Phi_{n}(0,X^{n}_{0})\|_{L^{m}(\Omega)}+1+\int_{0}^{t}\|Y^{n}_{s}\|_{L^{m}(\Omega)}{\mathord{{\rm d}}}s.$
By Gronwall’s inequality and the fact that $|X^{n}_{t}|=|\Phi_{n}^{-1}(t,Y^{n}_{t})|\lesssim 1+|Y^{n}_{t}|$, this implies (4.23).
(Step 2). Given $f\in C^{\infty}_{c}([0,T]\times{\mathbb{R}}^{d})$, let $u_{n}\in{\mathbb{L}}^{\infty}_{T}C^{\infty}_{b}$ be the unique smooth solution to the following backward PDE
$\partial_{t}u_{n}+\Delta u_{n}-\lambda u_{n}+b_{1,n}\cdot\nabla u_{n}+f=0,\ \ u_{n}(T)=0.$
For any $(\alpha,p,q)\in{\mathbf{I}}^{\alpha_{b}}_{p_{b},q_{b}}$, by (3.3) and (4.22), there is a $\lambda_{0}\geqslant 1$ so that for all $\theta\in(0,2+\alpha_{b}+\frac{d}{p_{b}}-\frac{2}{q_{b}}]$ and $\lambda\geqslant\lambda_{0}$,
By Itô’s formula, we have
$u_{n}(t,X^{n}_{t})=u_{n}(0,x)+\int^{t}_{0}(\lambda u_{n}-f)(s,X^{n}_{s}){\mathord{{\rm d}}}s+{\int_{0}^{t}(b_{2}\cdot\nabla u_{n})(s,X^{n}_{s}){\mathord{{\rm d}}}s}+\sqrt{2}\int^{t}_{0}\nabla u_{n}(s,X^{n}_{s}){\mathord{{\rm d}}}W_{s}.$
$\left|\int^{\tau_{1}}_{\tau_{0}}f(s,X^{n}_{s}){\mathord{{\rm d}}}s\right|\leqslant 2\|u_{n}\|_{\infty}+\lambda\delta\|u_{n}\|_{\infty}+{\|\nabla u_{n}\|_{\infty}\!\!\int_{\tau_{0}}^{\tau_{1}}|b_{2}|(s,X^{n}_{s}){\mathord{{\rm d}}}s}+\sqrt{2}\left|\int^{\tau_{1}}_{\tau_{0}}\nabla u_{n}(s,X^{n}_{s}){\mathord{{\rm d}}}W_{s}\right|.$
By BDG’s inequality and estimate (4.23), we have
$\left\|\int^{\tau_{1}}_{\tau_{0}}f(s,X^{n}_{s}){\mathord{{\rm d}}}s\right\|_{L^{m}(\Omega)}$ $\lesssim(1+\lambda\delta)\|u_{n}\|_{\infty}+\|\nabla u_{n}\|_{\infty}\delta\left\|\sup_{t\in[0,T]}(c_{0}+c_{1}|X^{n}_{t}|)\right\|_{L^{m}(\Omega)}$
$+m\left\|\left(\int^{\tau_{1}}_{\tau_{0}}|\nabla u_{n}(s,X^{n}_{s})|^{2}{\mathord{{\rm d}}}s\right)^{1/2}\right\|_{L^{m}(\Omega)}$
$\lesssim$ $(1+\lambda\delta)\|u_{n}\|_{\infty}+(m\delta^{1/2}+\delta(c_{0}+c_{1}\ell_{m}))\|\nabla u_{n}\|_{\infty}.$
by (4.24) and embedding (2.8), we have
and
Thus
$\left\|\int^{\tau_{1}}_{\tau_{0}}f(s,X^{n}_{s}){\mathord{{\rm d}}}s\right\|_{L^{m}(\Omega)}\lesssim$ $\left[(1+\lambda\delta)(1+\lambda)^{-\frac{1+\theta_{0}}{2}}+m\sqrt{\delta}(1+\lambda)^{-\frac{\theta_{0}}{2}}\right]\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q}}$
$+\delta(c_{0}+c_{1}\ell_{m})\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q}}$
In particular, if we take $\lambda=\lambda_{0}\vee\delta^{-1}$, then for all $\delta\in(0,1)$,
$\left\|\int^{\tau_{1}}_{\tau_{0}}f(s,X^{n}_{s}){\mathord{{\rm d}}}s\right\|_{L^{m}(\Omega)}\lesssim(m+c_{0}+c_{1}\ell_{m})\delta^{\frac{1+\theta_{0}}{2}}\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q}}.$ (4.25)
By a standard approximation, the above estimate still holds for any $f\in{\mathbb{L}}^{1}_{T}C_{b}\cap{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q}$. ∎
Remark 4.7.
By Stirling’s formula, when $c_{1}=0$, estimate (4.25) also implies that for some $c>0$,
$\sup_{n}{\mathbf{E}}\exp\left\{c\left|\int^{T}_{0}f(s,X^{n}_{s}){\mathord{{\rm d}}}s\right|/\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q}}\right\}<\infty.$
@@
Suppose that ($\widetilde{\bf H}^{\rm sub}_{b}$) holds.
  1. (i)
    For any $\mu\in{\mathcal{P}}({\mathbb{R}}^{d})$ having finite $m$-order moment, where $m$ satisfies (1.5), there exists a weak solution $(\mathfrak{F},X,W)$ to the SDE (1.1) with initial distribution $\mu$. Uniqueness holds within the class such that for any $(\alpha,p,q)\in{\mathbf{I}}^{\alpha_{b}}_{p_{b},q_{b}}$, $(X,{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q})$ satisfies Krylov’s estimate with parameters $(m,\theta_{0})$, where $\theta_{0}$ is as defined in (1.5).
  2. (ii)
    If $c_{1}=0$ in ($\widetilde{\bf H}^{\rm sub}_{b}$), then the moment assumption on $\mu$ can be dropped; and for any $t\in(0,T]$ and $X_{0}=x$, $X_{t}$ admits a density $\rho_{t}(x,x^{\prime})$ that satisfies the following two-sided Gaussian estimate: for all $x,x^{\prime}\in{\mathbb{R}}^{d}$,
    $C_{0}t^{-d/2}\mathrm{e}^{-\gamma_{0}|x-x^{\prime}|^{2}/t}\leqslant\rho_{t}(x,x^{\prime})\leqslant C_{1}t^{-d/2}\mathrm{e}^{-\gamma_{1}|x-x^{\prime}|^{2}/t},$ (4.26)
    where $C_{0},C_{1},\gamma_{0},\gamma_{1}>0$ depend only on the parameters $T,d,\alpha_{b},p_{b},q_{b}$, and $\|b\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}$.
@@
(Existence). Consider the approximation SDE (4.21). By Lemma 4.6, the law of $X^{n}$ in ${\mathbb{C}}_{T}$ is tight. Therefore, by Prohorov’s theorem, it is relatively compact in ${\mathcal{P}}({\mathbb{C}}_{T})$. Using the weak convergence method, it is standard to derive the existence of a weak solution (see [49] and also the proof of Theorem 5.7 below).
(Uniqueness and heat kernel estimate). By Lemma 4.5, since the transformed SDE (4.16) has a linear growth drift and bounded, uniformly non-degenerate Hölder continuous diffusion coefficients, it is well known that SDE (4.16) admits a unique weak solution $Y_{t}$ (see [50]). Consequently, weak uniqueness also holds for the original SDE (1.1).
When $c_{1}=0$, the Krylov constant in (4.3) does not depend on the initial distribution, allowing us to drop the moment assumption. Moreover, in this case, the transformed SDE (4.16) has a bounded drift. It is well known that for a starting point $Y_{0}=y$, the process $Y_{t}$ has a density $\tilde{\rho}_{t}(y,y^{\prime})$, which satisfies the following two-sided Gaussian estimate (for example, see [9]):
@@
The desired estimate (4.26) then follows by a change of variables (see [44]). ∎
Up to this point, we have assumed the condition $p_{b},q_{b}<\infty$ in ($\widetilde{\bf H}^{\rm sub}_{b}$). It must be noticed that for $q=\infty$ or $p=\infty$, the space $C_{b}({\mathbb{R}}^{d})$ is not dense in $L^{\infty}$, and consequently, the convergence property (4.2) does not holds. However, when $q=\infty$ or $p=\infty$, for $f\in{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,\infty}$, it follows from (2.4) that for any $\alpha^{\prime}<\alpha$,
$\|f_{n}-f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha^{\prime}}_{p,\infty}}\lesssim\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,\infty}}n^{\alpha^{\prime}-\alpha}.$ (4.27)
This provides an alternative form of convergence even in cases where $q=\infty$ or $p=\infty$.
@@
Since $b_{1},\mathord{{\rm div}}b_{1}\in{\mathbb{L}}^{q_{b}}_{T}{\mathbf{H}}^{\alpha_{b}}_{p_{b}}$, noting that $\alpha_{b}>-2-\alpha_{b}$, by Lemma 2.4,
@@
one sees that ($\widetilde{\bf H}^{\rm sub}_{b}$) holds. If $q_{b},p_{b}<\infty$, then the desired results follow by the above theorem.
For the case $q_{b}=\infty$ or $p_{b}=\infty$, since $\frac{2}{q_{b}}+\frac{d}{p_{b}}<1+\alpha_{b}$, one can choose $\tilde{\alpha}_{b}\in(-1,\alpha_{b})$ and $\tilde{q}_{b}<q_{b}\leqslant\infty$ so that $\frac{2}{\tilde{q}_{b}}+\frac{d}{p_{b}}<1+\tilde{\alpha}_{b}$. In particular, by (2.8) we have
$\kappa_{b}=\|b_{1}\|_{{\mathbb{L}}^{\tilde{q}_{b}}_{T}{\mathbf{B}}^{\tilde{\alpha}_{b}}_{p_{b},\tilde{q}_{b}}}\leqslant\|b_{1}\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}<\infty,$
and
$\tilde{\kappa}_{b}=\|\mathord{{\rm div}}b_{1}\|_{{\mathbb{L}}^{\tilde{q}_{b}}_{T}{\mathbf{B}}^{-2-\tilde{\alpha}_{b}}_{p_{b},1}}\leqslant\|\mathord{{\rm div}}b_{1}\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}<\infty,$
and by (4.27),
$\|b_{1,n}-b_{1}\|_{{\mathbb{L}}^{\tilde{q}_{b}}_{T}{\mathbf{B}}^{\tilde{\alpha}_{b}}_{p_{b},\tilde{q}_{b}}}\leqslant\|b_{1,n}-b_{1}\|_{{\mathbb{L}}^{\tilde{q}_{b}}_{T}{\mathbf{B}}^{(\alpha_{b}+\tilde{\alpha}_{b})/2}_{p_{b},\infty}}\leqslant C\|b_{1}\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}n^{(\tilde{\alpha}_{b}-\alpha_{b})/2},$
and similarly,
$\lim_{n\to\infty}\|\mathord{{\rm div}}b_{1,n}-\mathord{{\rm div}}b_{1}\|_{{\mathbb{L}}^{\tilde{q}_{b}}_{T}{\mathbf{B}}^{-2-\tilde{\alpha}_{b}}_{p_{b},1}}=0.$
Thus one can check that Lemma 4.5 still holds. The rest is the same as Theorem 4.8. ∎